OpenStudy (anonymous):

For a sine function with amplitude =0.75 and period =10 , what is y(4) ?

OpenStudy (anonymous):

Depends on the initial angle.

OpenStudy (anonymous):

There's no phase shift mentioned, so I think we can assume it starts at x=0.

OpenStudy (anonymous):

You need to develop a function of the form \[\large y=Asin(Bx+C)\] A=amplitude B=frequency (for sine, it is 2π/T, T=period) C=phase shift

OpenStudy (anonymous):

The function sin x has amplitude 1 and period \[2\pi\] Therefore your function can be expressed as \[0.75\sin \frac{ 2\pi }{ 10 }x\] Just stick that with the right numbers into a calculator and you're sorted!